# MSACL: Multi-Step Actor-Critic Learning with Lyapunov Certificates for Exponentially Stabilizing Control

### v1

@@ VI-A Discussion
This subsection provides an analytical discussion on the functional roles of the various components within the MSACL framework. As highlighted in Remark 4, MSACL extends the Soft Actor-Critic (SAC) algorithm by integrating multi-step Lyapunov certificate learning, thereby establishing formal stability guarantees while preserving the exploration efficiency of maximum entropy reinforcement learning. In model-free stabilization and tracking tasks, conventional SAC, which lacks the structural guidance provided by Lyapunov theory, may fail to converge to stable policies under improper reward structures. Even with well-designed rewards, standard RL often requires exhaustive exploration of the action space to identify stabilizing directions, leading to prohibitive training times. To address these limitations, our approach introduces a Lyapunov network that serves as a stability-oriented critic, providing explicit directional guidance during policy updates by identifying action sequences that facilitate the descent of the Lyapunov certificate.
The theoretical foundation of this mechanism is rooted in Lemma 1, which establishes the equivalence between the Lyapunov stability conditions and exponential state convergence. To ensure the Lyapunov certificate $V_{\psi}$ is learned accurately, we developed a training methodology that aligns theoretical requirements with empirical trajectories. Specifically, the introduction of Exponential Stability Labels (ESL) in Definition 2 allows the algorithm to categorize trajectory segments as positive or negative samples based on their adherence to exponential decay criteria. This classification enables the construction of the loss function (11), which effectively bridges abstract stability conditions with observed system dynamics.
The policy optimization process leverages this learned certificate through the stability advantage mechanism defined in (17). This advantage function quantifies the rate of Lyapunov descent along trajectories, prioritizing policy updates that simultaneously ensure stability and accelerate convergence. Our experimental results confirm that this dual-objective optimization yields policies that achieve both stability and rapid convergence across diverse and challenging benchmarks. A critical insight is that the efficacy of the policy is fundamentally tied to the fidelity of the learned Lyapunov certificate. The multi-step sampling strategy, combined with the ESL-based loss formulation, ensures that $V_{\psi}$ accurately reflects the theoretical stability criteria, providing reliable gradients for policy improvement.
While developed for stabilization and tracking, the MSACL framework is highly extensible to other stability-critical applications. Potential extensions include trajectory planning for autonomous vehicles [42], spacecraft proximity operations [43], and robotic manipulation tasks requiring rigorous safety guarantees [44, 45]. The integration of Lyapunov-based guarantees within maximum entropy RL represents a significant step toward developing verifiable, learning-based control systems for real-world deployment.

### v2

@@ VI-A Discussion
This subsection elucidates the functional roles of MSACL components. As highlighted in Remark 4, MSACL augments the SAC framework with a stability-oriented critic, departing from standard RL that relies on exhaustive exploration to discover stabilizing behaviors. Instead, MSACL leverages a learned Lyapunov certificate to provide explicit, structured guidance for policy updates.
Theoretically, MSACL is grounded in Lemma 1, which establishes the equivalence between Lyapunov stability and exponential convergence. To operationalize this exponential stability requirement, the ESL mechanism (Definition 2) categorizes trajectory segments based on their contraction properties, transforming the stability condition into a supervised learning task. This ensures the synthesis of a high-fidelity Lyapunov certificate $V_{\psi}$ that accurately captures the underlying system dynamics. Such fidelity is crucial for providing the actor with theoretically sound, stability-aware gradients. Consequently, the stability advantage in (18) prioritizes actions that facilitate Lyapunov descent, effectively guiding the policy toward the region of stability.
While developed for stabilization and tracking, MSACL is extensible to safety-critical applications including autonomous vehicle trajectory planning [52], spacecraft proximity operations [53], and safety-guaranteed robotic manipulation [54, 55]. The integration of Lyapunov guarantees within MERL represents a step toward verifiable learning-based control.